OpenStudy (errrwut):

Need help understanding how an answer was achieved: (my worked out answer is inside) Particle physicists use particle track detectors to determine the lifetime of short-lived particles. A muon has a mean lifetime of 2.2 micro/seconds and makes a track 9.5 cm long before decaying into an electron and two neutrinos. What was the speed of the muon?

OpenStudy (errrwut):

\[T_0 = 2.2 µs\] \[L = 9.5cm\] Using, \[\beta = \frac{L*\sqrt{1-\beta^2}}{c*T_0}\] Solving for \[\beta\] \[\beta=\frac{L}{\sqrt{L^2+c^2*T^2}} \] You can neglect c for the mean time? I get \[\beta = 0.974218\] However, the solution is supposedly \[\beta = 1.4*10^{-4}\]

OpenStudy (osprey):

Question reminds me of the "Mt Washington expt/demo" done by MIT and printed in a book by AP FRENCH. I've had a look at this and am confused. You want the SPEED of the muon, but you quote an "answer" of 0.0001 no units for something (beta according to the post). I doubt that that's the SPEED of the muon, 'cos it's going fairly fast. An approximation based on LOW SPEED mechanics gives me 40, 000 m/s for the speed. This is well below c, and MAY make this a "classical" problem. Maybe the muon speed is much higher than that approx, but its a lot bigger for a SPEED numbe than is the "beta value" in the post. http://perendis.webs.com

OpenStudy (errrwut):

Yeah, \[\beta = \frac{v}{c}\] Solving for v will get you \[v = 1.4*10^{-4}*c\]

OpenStudy (osprey):

At the END of the muon's journey it may very well be going at the lower speed. Problem seems to be to work backwards to what "caused" that particular speed. Being a high speed muon, special rel is implied ?

OpenStudy (errrwut):

I'll show you the solution from the text book: In the muon’s frame \[T_0 =  2.2 μs \] In the lab frame the time is longer; see Equation (2.19)*: \[T' = T_0.\] In the lab the distance traveled is \[9.5cm = v*T' = v * \gamma * T_0 = \beta * c * \gamma * T_0 \] since \[v = \beta c.\] Therefore, \[\beta = \frac{9.5cm(\sqrt{1- \beta ^2})}{cT_0}, so\] \[\beta = \frac{v}{c} =\frac{9.5cm(\sqrt{1- \beta ^2})}{c (2.2\mu s)}.\] Now all quantities are known except \[\beta .\] Solving for \[\beta \] we find \[\beta = 1.4*10^{-4}\] or \[ v = 1.4*10^{-4}c\] *For more context this is how Equation 2.19 is defined in the text book: \[T' = \frac{T_0}{\sqrt{1 - \frac{v^2}{c^2}}} = \gamma*T_0\]

OpenStudy (errrwut):

Correction to line in the above comment: "In the lab frame the time is longer; see Equation (2.19)*: \[T' = T_0\]" The equation is incorrect I forgot to type the gamma. It should be: In the lab frame the time is longer; see Equation (2.19)*: \[T' = \gamma * T_0\]